OpenStudy (anonymous):

Nick, a mattress salesman, earns a base salary of $2,000 per month and a 5% commission on all of his sales. If he earned $2,350 in June, how much did he sell?

jimthompson5910 (jim_thompson5910):

if s = amount of sales (in dollars), then E = 2000 + 0.05s where E is the amount Nick earns

jimthompson5910 (jim_thompson5910):

So "If he earned $2,350 in June", then E = 2350

jimthompson5910 (jim_thompson5910):

which means we can say E = 2000 + 0.05s 2350 = 2000 + 0.05s

jimthompson5910 (jim_thompson5910):

Keep going to solve for 's'